OpenStudy (anonymous):

1kg is equal to 100 g 10000 dg 1000 mg 1000 cg

OpenStudy (anonymous):

100 g lol

OpenStudy (anonymous):

1 kg = 1000 gm

OpenStudy (anonymous):

10000 dg (1dg = 1/10g; 1g = 1/1000kg)

OpenStudy (anonymous):

confushed to many different answers

OpenStudy (anonymous):

1 kg = 1000 g = 10000 dg = 1000000mg so its 10000 dg

OpenStudy (anonymous):

you always have to trust the engineer lol

OpenStudy (anonymous):

1 Gram = 1000 Milligram 1 Decigram = 10 centigram 1 Decagram = 1000 centigram I'll write the table see that.. 10 milligram = 1 centigram 10 centigram = 1 decigram 10 decigram = 1 gram 10 gram = 1 decagram

OpenStudy (anonymous):

This may help

OpenStudy (anonymous):

@surdawi isnt that the truth :)

OpenStudy (anonymous):

thanks @surdawi